Dantzler Corporation is a fast-growing supplier of office products. Analysts project the following free cash flows (FCFs) during the next 3 years, after which FCF is expected to grow at a constant 5% rate. Dantzler's WACC is 10%.

Year 0 1 2 3
....... ....... ....... ....... ....... ....... ....... .......
....... ....... ....... ....... ....... ....... ....... ......
FCF ($ millions) - $25 $10 $47
  1. What is Dantzler's horizon, or continuing, value? (Hint: Find the value of all free cash flows beyond Year 3 discounted back to Year 3.) Enter your answer in millions. For example, an answer of $13,550,000 should be entered as 13.55. Do not round intermediate calculations. Round your answer to two decimal places.
    $   million
  2. What is the firm's market value today? Assume that Dantzler has zero non-operating assets. Enter your answer in millions. For example, an answer of $13,550,000 should be entered as 13.55. Do not round intermediate calculations. Round your answer to two decimal places.
    $   million
  3. Suppose Dantzler has $32.70 million of debt and 30 million shares of stock outstanding. What is your estimate of the current price per share? Write out your answer completely. For example, 0.00025 million should be entered as 250. Do not round intermediate calculations. Round your answer to the nearest cent.
    $  

Grade it Now

Save & Continue
Answer #1

a. Horizon value at Year3=FCF4/(WACC-growth rate)

FCF4=FCF3*(1+growth rate)=$47*(1+5%)=$49.35 million

Horizon value=$49.35/(10%-5%)=$987.0 million

b. Value of the Firm=(FCF1/(1+10%))+(FCF2/(1+10%)^2)+((FCF3+Horizon value at year3/(1+10%)^3)

=(-25/1.10)+(10/1.10^2)+(1034/1.10^3)

=-22.73+8.26+776.86

=$762.40 million

c. Value of the equity=value of the firm-value of debt=$762.40-$32.7=$729.70 million

Value of share=Value of the equity/Shares outstanding=729.70/30=$24.32